# Text to model via SysML: Automated generation of dynamical system computational models from unstructured natural language text via enhanced System Modeling Language diagrams

### v1

@@ Abstract
This paper contributes to speeding up the design and deployment of engineering dynamical systems by proposing a strategy for exploiting domain and expert knowledge for the automated generation of dynamical system computational model starting from a corpus of document relevant to the dynamical system of interest and an input document describing the specific system. This strategy is implemented in five steps and, crucially, it uses system modeling language diagrams (SysML) to extract accurate information about the dependencies, attributes, and operations of components. Natural Language Processing (NLP) strategies and Large Language Models (LLMs) are employed in specific tasks to improve intermediate outputs of the SySML diagrams automated generation, such as: list of key nouns; list of extracted relationships; list of key phrases and key relationships; block attribute values; block relationships; and BDD diagram generation. The applicability of automated SysML diagram generation is illustrated with different case studies. The computational models of complex dynamical systems from SysML diagrams are then obtained via code generation and computational model generation steps. In the code generation step, NLP strategies are used for summarization, while LLMs are used for validation only. The proposed approach is not limited to a specific system, domain, or computational software. The applicability of the proposed approach is shown via an end-to-end example from text to model of a simple pendulum, showing improved performance compared to results yielded by LLMs only.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have received particular attention for code completion, synthesis, analysis, and interpretation of the results, and by converting natural language text into domain-specific languages for interaction with software tools [7, 8, 9, 10, 11]. The use of LLMs for generating dynamic models of chemical reaction systems from descriptions written in natural language has recently been explored [11]. Specifically, the Llama 3.1- 8B LLM model [12] was fine-tuned an LLM using low-rank adaptation on synthetically generated Modelica code for different reactor scenarios to increase the semantic and syntactic accuracy of automatically generated domain-specific programming language instructions in Modelica [11]. However, these LLMs-based approaches are often limited in their ability to generalise to unseen dynamical system simulations scenarios that were not included in the training data, as demonstrated and discussed in [11]. Moreover, these approaches do not provide a structured representation of the system, which is crucial for understanding the interaction between components.
@@
The proposed approach does not intend to replace engineering personnel. Instead, it is intended to be used to assist engineers in the generation of computational models for dynamical systems from unstructured natural language text. The proposed approach improves performance with respect to LLMs, and it yields intermediate outputs that can be controlled by the engineering team. Consequently, it would enable faster investigations of the dynamical system performance by investigating simulations of the system under various operational and environmental conditions, including extreme events. The proposed approach is designed to be general and applicable to a wide range of dynamical systems, allowing for the generation of computational models from unstructured natural language text only. The approach is not limited to any specific type of system or domain (i.e., it is not context specific), making it versatile and adaptable to various engineering contexts and in other scientific fields. Moreover, each step in this approach is not implementation-specific, meaning that each step could be carried out by exploiting alternative tools or algorithms that produce the same outputs for the set of given inputs at a step. This allows future implementations of the approach to be carried out using improved tools that outperform those that exist today or are more suited to the task at hand. The key enabler of the proposed approach is exploiting and enhancing strategies for automatically generating system representations in the form of SysML diagrams [6, 22] as an intermediate step towards generating computational models. The code generation step converts the SysML diagram into code that models the current dynamical system behavior and can be modified by the user for further investigations. Finally, the computational model generation step uses the generated code to run simulations of the dynamical system across various input conditions.
@@
    The design of an automated approach to generate dynamical system computational models from natural language text by enhancing SysML.
@@
The proposed approach enables the automated generation of computational models to investigate the performance of a dynamical system starting from domain and expert knowledge in the form of unstructured natural language text describing the dynamical system of interest by exploiting System modeling Language (SysML) diagrams, and it consists of the steps summarised in Figure 1.
@@
      Function Body Generation: The code skeleton and code comments are used to implement the function bodies with the relevant equations of motion in the class skeleton.
@@ 3 Context-free automated generation of SysML diagrams from unstructured natural language text
The procedure for automatic generation of the SysML diagram from unstructured natural language text builds on the strategy proposed in [6] and consists of the following steps:
@@ 3.1 Text preparation and pre-processing
The unimodal data source used within the proposed approach is unstructured natural language text without a predefined writing style, describing the dynamical system for which a computational model is going to be automatically built. A corpus of text documents describing the class of dynamical systems of interest is manually selected and used to automatically extract information to obtain a computational model of a specific dynamical system of interest to the user. Information contained in tables, equations, and figures included in each document is discarded, and only textual information is retained. Based on the results obtained in [6] for automated generation of SysML, the size of the corpus could be considered sufficient when the document count is larger than 100, and the average word count per document is larger than 500. Typically, this size of corpus is obtained by splitting large documents (e.g., books, manuals, reports, patents, standards) into chapters and sections to achieve a sufficient document count [6]. However, when dealing with documentation referring to a specific dynamical system, it might not be possible to have such large corpus of documents. Therefore, in the present work a range of examples is considered. In the first case, the corpus consist of the chapters of the book [24], and therefore comprises less than 30 documents. In the second case, a corpus of patents with only 10 documents is considered. The third example, that is used as an end-to-end case study in Section 6 consider a document under 500 words and uses the same corpus of patents. A summary of the input documents and corpuses used in the paper is shown in Table 8.
@@ 4 Steps for implementation and validation of context-free automated generation of SysML diagrams
The steps for the automatic generation of the BDD diagrams to enable the automated evaluation of a dynamical system computational model can be summarized as follow:
@@ 4.1.1 Accuracy of key phrase extraction
The validation results of the key phrase extraction are shown in Figure 5.
The results show that the current methodology outperforms GPT-4o [^1] The prompts for this can be found at https://chatgpt.com/share/67771274-b768-800f-9ef4-755172c091de in key phrase extraction on recall (fraction of ground truth phrases extracted) whilst performing slightly worse on precision (fraction of extracted phrases that are in the ground truth). The results also show that the preprocessing improvements marginally improve the key phrase extraction performance, more notable in terms of recall.
@@ 4.1.2 Observations on the information encoded in the BDD diagrams
The first example is focused on a patent document on the ventilation of internal combustion engines [45], where information are expressed in context-rich images and in text. This is particularly challenging for the proposed approach since only text is retained and images are discarded. The patent document [45] whose excerpt can be found in Table 8 was used, together with a corpus of 10 patents. The generated BDD sub-diagram under the “turbine” block is shown in Figure 19. Here the information extracted from the generated BDD diagram is directly compared with the source text (The comparison is done on extracts where the text corresponds to the components in the sub-diagram as the sub-diagram is not a complete representation of the entire system described by the patent). It can be observed that the BDD diagram successfully extracts meaningful components of the system as blocks, as components such as “turbine” and “turbine housing” which are components mentioned in the source text were extracted. However, due to the many numbers in the text indicating the components in the figures of the patent document (e.g. “The exhaust valve cam 46a is in contact with the end of the exhaust valve 40,”), the BDD diagram incorrectly identifies some of them as attributes. This is due to the inability of the current methodology to extract information from the context-rich images in the patent. The numerical values in the source text are, therefore, incorrectly identified as attribute values for the components of the system. This example supports the use of multimodal information extraction for generating the BDD diagrams, which is beyond the scope of the present work. In any case, in the context of engineering practices, it might be time consuming and error prone, to check the information encoded in the entire BDD diagram generated with respect to the source document.
@@ 4.1.3 Accuracy of blocks and attributes from the extracted BDD Diagram
Even though there are well established methods for evaluating attribute value extraction in its general case such as directly evaluating the precision, recall and F1 score by directly comparing attributes extracted from the same sentence [46]. To the best of the authors’ knowledge, the assessment of the accuracy of the blocks and attribute value extraction of multiple components in dynamical systems has not been addressed in the literature.
@@
The low rate of correctly identified attributes and bad matching of block labels could be due to the older language in the document. Hence, a more modern sample text describing a hydraulic robotic arm system was generated using GPT-4o and investigated the performance of the attribute value extraction on this text. And excerpt of the sample text is shown in Figure 7. The results of the attribute value extraction on this text are shown in Table 2. The results show that the attribute value extraction is slightly more accurate than the previous text, marginally confirming our hypothesis that it is easier to extract attributes of components from modern text. The zero attribute match score for this text is 1 mainly because all the ground truth blocks have attributes in them. The normalized set match score was much higher than the previous text, showing that the extracted block labels match very well with the ground truth block labels. Even though overall the result scores are not extremely high, this is still a significant improvement over the algorithm without this step as at least some attributes of the system components are extracted.
@@
The performance of the attribute value extraction on a synthetic text describing a hydraulic system was also assessed without the improved preprocessing steps. The results are shown in Table 3. The results show that the attribute value extraction marginally less accurate than the previous example. However, the zero attribute match score and normalized set match score stayed the same. This demonstrates the importance of improved preprocessing steps in the attribute value extraction process.
@@ 5 From SysML diagrams to dynamical system computational models
The task of creating computational models from SysML diagrams is typically carried out manually by using XMI (XML Metadata Interchange) to add relevant metadata (such as equations) from a database of relevant information into the SysML to form a meta-model. The meta-model is then parsed to generate a script in the target simulation language [17]. However, it is currently not possible to generally generate the meta-model for general dynamical systems.
@@ 5.1.1 File Structure Generation
This step concerns the generation of a file structure that helps organize the generated code for each component the computational model. A good file structure is one that mimics the structure of the dynamical system itself.
The dependencies of components can be extracted from the composition relationships in the BDD diagram. Specifically, a block that is a part of a parent block should have its code generated in a file living in a subdirectory of the file representing the parent block that it is a part of. This approach generates a folder hierarchy with blank files inside them to be populated in further steps.
The following is the sequence of steps are automated to generate the file structure from a BDD:
@@ 5.1.2 Class Skeleton Generation
    Thee import statements required for the class to work are automatically added at the the top of the file. This was done automatically by deterministically determining which classes need to be imported by inspecting the generalization and composite relationships tied to the block being considered. For each generalization relationship, the file representing the block at the specific end of the relationship should import its general parent block’s file from its file living on the same folder level. For each composite relationship, the file representing the block on the composite end of the relationship should import the child block’s file living in a subfolder level relative to the current file. More imports can be included, as described in Section 5.2.
@@ 5.1.3 Code Comment Generation
    The Pegasus extraction model [52] was used to generate the abstractive summaries with different minimum lengths specified. The Pegasus model is a transformer-based model that is trained on a large corpus of text to generate abstractive summaries. The model works by generating a sequence of words that captures the main ideas of the input text. The results of using the Pegasus model to generate the abstractive summaries are shown in Figure 9. It is shown that even when the minimum length is set to 25, the model hallucinates information that is not present in the input text (there was no mention about the importance of the rudder in the input text), and that is not relevant for the modeller. This is a common problem with abstractive summarization models and is one of the reasons why it is better to use extractive summarization for our task. This problem only worsens when the minimum length is increased to 60, where the model generates a summary that is not even close to the input text.
@@ 5.2.1 Function Body Generation strategy
    A list of dynamical equation implementation templates to choose from, an informal description of the system, the states of the equations, and the necessary parameters for the equation are shown in Table 10, with an example implementation template written in Jinja2 (a templating language) shown in LABEL:lst:example_dynamical_equation_template. This has to be manually compiled by engineers with domain knowledge for the approach.
@@ 5.3 Computational model generation
When having a working code that models each of the separate components of the dynamical system, a computational model is obtained by combining these components to produce a simulation of the system is necessary to draw conclusions about the performance of the system. This can be done by creating a highest level system class that initialises the class instances of all the components of the system, sets all the initial conditions and simulates each of them at each timestep. An example system file can be seen in LABEL:lst:example_system_class. The simulation results (i.e., the values of the states of the dynamical system) can then be used to compute performance metrics for the system or to verify whether the states exceed safe values for a system’s operation.
The generation of the system class is done by initialising all of the top-level components that subsequently initialise all their child components in an initialization function (the function called during the construction of a class instance). A class function named “simulate” is then created for the system class. This “simulate” function takes in simulation parameters and initial conditions as user input. Simulation parameters would include the length of each timestep and the total time of the simulation. The required initial conditions for user input depend on the dynamical equation being simulated. The outputs of these simualte functions will be the states of each component in the dynamical system at each timestep. These are be compiled into arrays of states at each timestep forming the output of the simulation.
@@ 6.3 Computational model generation
7 Discussions
The results presented demonstrate that the proposed approach is capable of automatically generating the computational model of a dynamical system directly from unstructured natural language text by exploiting System modeling Language (SysML) diagrams and by making strategic use of NLP strategies and very limited use of LLMs.
By investigating BDD diagram results for different input texts, it was shown the advantage of the proposed approach over zero-shot LLM prompting. It was demonstrated that the generated BDD diagrams can sufficiently accurately represent the structure and components of dynamical systems and their relations. Furthermore, new validation strategies for blocks with attributes were proposed, providing a foundation for future research in this area and enabling the engineer to identify steps which might require modifications.

### v2

@@ Abstract
This paper contributes to speeding up the design and deployment of engineering dynamical systems by proposing a strategy for exploiting domain and expert knowledge for the automated generation of a dynamical system computational model starting from a corpus of documents relevant to the dynamical system of interest and an input document describing the specific system. This strategy is implemented in five steps and, crucially, it uses system modeling language diagrams (SysML) to extract accurate information about the dependencies, attributes, and operations of components. Natural Language Processing (NLP) strategies and Large Language Models (LLMs) are employed in specific tasks to improve intermediate outputs of the SySML diagrams automated generation, such as: list of key nouns; list of extracted relationships; list of key phrases and key relationships; block attribute values; block relationships; and BDD diagram generation. The applicability of automated SysML diagram generation is illustrated with different case studies. The computational models of complex dynamical systems from SysML diagrams are then obtained via code generation and computational model generation steps. In the code generation step, NLP strategies are used for summarization, while LLMs are used for validation only. The proposed approach is not limited to a specific system, domain, or computational software. Domain and expert knowledge is integrated by providing a set of equation implementation templates. The applicability of the proposed approach is shown via an end-to-end example from text to model of a simple pendulum, showing improved performance compared to results yielded by LLMs only.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have received particular attention for code completion, synthesis, analysis, and interpretation of the results, and by converting natural language text into domain-specific languages for interaction with software tools [7, 8, 9, 10, 11]. The use of LLMs for generating dynamic models of chemical reaction systems from descriptions written in natural language has recently been explored [11]. Specifically, the Llama 3.1- 8B LLM model [12] was fine-tuned by using low-rank adaptation on synthetically generated Modelica code for different reactor scenarios to increase the semantic and syntactic accuracy of automatically generated domain-specific programming language instructions in Modelica [11]. However, these LLMs-based approaches are often limited in their ability to generalise to unseen dynamical system simulations scenarios that were not included in the training data, as demonstrated and discussed in [11]. Moreover, these approaches do not provide a structured representation of the system, which is crucial for understanding the interaction between components.
@@
The proposed approach does not intend to replace engineering personnel. Instead, it is intended to be used to assist engineers in the generation of computational models for dynamical systems from unstructured natural language text by exploiting domain and expert knowledge. The proposed approach improves performance with respect to LLMs, and it yields intermediate outputs that can be controlled by the engineering team. Consequently, it would enable faster investigations of the dynamical system performance by investigating simulations of the system under various operational and environmental conditions, including extreme events. The proposed approach is designed to be general and applicable to a wide range of dynamical systems, allowing for the generation of computational models from unstructured natural language text only. The approach is not limited to any specific type of system or domain (i.e., it is not context specific), making it versatile and adaptable to various engineering contexts and in other scientific fields. Moreover, each step in this approach is not implementation-specific, meaning that each step could be carried out by exploiting alternative tools or algorithms that produce the same outputs for the set of given inputs at a step. This allows future implementations of the approach to be carried out using improved tools that outperform those that exist today or are more suited to the task at hand. The key enabler of the proposed approach is exploiting and enhancing strategies for automatically generating system representations in the form of SysML diagrams [6, 22] as an intermediate step towards generating computational models. The code generation step converts the SysML diagram into code that models the current dynamical system behavior and can be modified by the user for further investigations. Domain and expert knowledge on the specific dynamical system is integrated in this step by including manually compiled templates of equation implementation. Finally, the computational model generation step uses the generated code to run simulations of the dynamical system across various input conditions.
@@
    An automated approach to generate dynamical system computational models from natural language text by enhancing SysML.
@@
The proposed approach enables the automated generation of computational models to investigate the performance of a dynamical system starting from domain and expert knowledge in the form of unstructured natural language text describing the dynamical system of interest by exploiting System modeling Language (SysML) diagrams, and in the form of a set of code templates. It consists of the steps summarised in Figure 1.
@@
      Function Body Generation: The code skeleton and code comments are used to implement the function bodies with the relevant equations of motion in the class skeleton. Within this step, a set of equation implementation templates is provided by engineers with domain knowledge on the system.
@@ 3 Context-free automated generation of SysML diagrams from unstructured natural language text
The procedure for automatic generation of the SysML diagram from unstructured natural language text. The approach is not tuned to a specific system or domain (i.e., it is context-free), and it builds on the strategy proposed in [6]. It consists of the following steps:
@@ 3.1 Text preparation and pre-processing
The unimodal data source used within the proposed approach is unstructured natural language text without a predefined writing style, describing the dynamical system for which a computational model is going to be automatically built. A corpus of text documents describing the class of dynamical systems of interest is manually selected and used to automatically extract information to obtain a computational model of a specific dynamical system of interest to the user. It is worth emphasizing here that information contained in tables, equations, and figures included in each document is discarded, and only textual information is retained. Based on the results obtained in [6] for automated generation of SysML, the size of the corpus could be considered sufficient when the document count is larger than 100, and the average word count per document is larger than 500. Typically, this size of corpus is obtained by splitting large documents (e.g., books, manuals, reports, patents, standards) into chapters and sections to achieve a sufficient document count [6]. However, when dealing with documentation referring to a specific dynamical system, it might not be possible to have such large corpus of documents. Therefore, in the present work a range of examples is considered. In the first case, the corpus consist of the chapters of the book [24], and therefore comprises less than 30 documents. In the second case, a corpus of patents with only 10 documents is considered. The third example, that is used as an end-to-end case study in Section 6 consider a document under 500 words and uses the same corpus of patents. A summary of the input documents and corpuses used in the paper is shown in Table 8.
@@ 4 Steps for implementation and validation of context-free automated generation of SysML diagrams
The steps for the automatic generation of the BDD diagrams to enable the automated evaluation of a dynamical system computational model can be summarized as follows:
@@ 4.1.1 Accuracy of key phrase extraction
The validation results of the key phrase extraction are shown in Figure 5. The results show that the current methodology outperforms GPT-4o [^1] The prompts for this can be found at https://chatgpt.com/share/67771274-b768-800f-9ef4-755172c091de in key phrase extraction on recall (fraction of ground truth phrases extracted) whilst performing slightly worse on precision (fraction of extracted phrases that are in the ground truth). The results also show that the preprocessing improvements marginally improve the key phrase extraction performance, more notable in terms of recall.
@@ 4.1.2 Observations on the information encoded in the BDD diagrams
The first example is focused on a patent document on the ventilation of internal combustion engines [45], where information are expressed in context-rich images and in text. This is particularly challenging for the proposed approach since only text is retained and images are discarded. The patent document [45] whose excerpt can be found in Table 8 was used, together with a corpus of 10 patents. The generated BDD sub-diagram under the “turbine” block is shown in Figure 19. Here the information extracted from the generated BDD diagram is directly compared with the source text (The comparison is done on extracts where the text corresponds to the components in the sub-diagram as the sub-diagram is not a complete representation of the entire system described by the patent). It can be observed that the BDD diagram successfully extracts meaningful components of the system as blocks, as components such as “turbine” and “turbine housing” which are components mentioned in the source text were extracted. However, due to the many numbers in the text indicating the components in the figures of the patent document (e.g. “The exhaust valve cam 46a is in contact with the end of the exhaust valve 40,”), the BDD diagram incorrectly identifies some of them as attributes. This is due to the inability of the current methodology to extract information from the context-rich images in the patent. The numerical values in the source text are, therefore, incorrectly identified as attribute values for the components of the system In the context of engineering practices, it might be time consuming and error prone, to check the information encoded in the entire BDD diagram generated by going through the entire source document. This example supports the use of multimodal information extraction for generating the BDD diagrams, which is beyond the scope of the present work.
@@ 4.1.3 Accuracy of blocks and attributes from the extracted BDD Diagram
The attribute value extraction is typically assessed with the precision, recall and F1 score by directly comparing attributes extracted from the same sentence [46]. To the best of the authors’ knowledge, the assessment of the accuracy of the blocks and attribute value extraction of multiple components in dynamical systems has not been addressed in the literature.
@@
The low rate of correctly identified attributes and bad matching of block labels could be due to the older language in the document. Hence, a more modern sample text describing a hydraulic robotic arm system was generated using GPT-4o and investigated the performance of the attribute value extraction on this text. And excerpt of the sample text is shown in Figure 7. The results of the attribute value extraction on this text are shown in Table 2. The results show that the attribute value extraction is slightly more accurate than the previous text, marginally confirming our hypothesis that it is easier to extract attributes of components from modern text. The zero attribute match score for this text is 1 mainly because all the ground truth blocks have attributes in them. The normalized set match score was much higher than the previous text, showing that the extracted block labels match very well with the ground truth block labels. Overall, the scores indicate an improvement with respect to the algorithm without this step as at least some attributes of the system components are extracted.
@@
The performance of the attribute value extraction on a synthetic text describing a hydraulic system was also assessed without the improved preprocessing steps. The results are shown in Table 3. The results show that the attribute value extraction is marginally less accurate than the previous example. However, the zero attribute match score and normalized set match score stayed the same. This demonstrates the importance of improved preprocessing steps in the attribute value extraction process.
@@ 5 From SysML diagrams to dynamical system computational models
The task of creating computational models from SysML diagrams is typically carried out manually by using XMI (XML Metadata Interchange) to add relevant metadata (such as equations) from a database of relevant information into the SysML to form a meta-model. The meta-model is then parsed to generate a script in the target simulation language [17]. To the best of the authors’ knowledge, it is currently not possible to automatically generate the meta-model of a general dynamical system from the SysML diagram.
@@ 5.1.1 File Structure Generation
This step concerns the generation of a file structure that helps organize the generated code for each component the computational model. A good file structure is one that mimics the structure of the dynamical system itself. The dependencies of components can be extracted from the composition relationships in the BDD diagram. Specifically, a block that is a part of a parent block should have its code generated in a file living in a subdirectory of the file representing the parent block that it is a part of. This approach generates a folder hierarchy with blank files inside them to be populated in further steps. The following is the sequence of steps are automated to generate the file structure from a BDD:
@@ 5.1.2 Class Skeleton Generation
    Thee import statements required for the class to work are automatically added at the the top of the file. This was done automatically by determining deterministically which classes need to be imported by inspecting the generalization and composite relationships tied to the block being considered. For each generalization relationship, the file representing the block at the specific end of the relationship should import its general parent block’s file from its file living on the same folder level. For each composite relationship, the file representing the block on the composite end of the relationship should import the child block’s file living in a subfolder level relative to the current file. More imports can be included, as described in Section 5.2.
@@ 5.1.3 Code Comment Generation
    The Pegasus extraction model [52] was used to generate the abstractive summaries with different minimum lengths specified. The Pegasus model is a transformer-based model that is trained on a large corpus of text to generate abstractive summaries. The model works by generating a sequence of words that captures the main ideas of the input text. The results of using the Pegasus model to generate the abstractive summaries are shown in Figure 9. It is shown that, as anticipated, the model will hallucinate information if given the task to produce a summary longer than the original text (up to 60 words). However, even when the minimum length is set to 25, the model hallucinates information that is not present in the input text (there was no mention about the importance of the rudder in the input text), and that is not relevant for the modeler. This is a common problem with abstractive summarization models and is one of the reasons why it is better to use extractive summarization for our task. This problem only worsens when the minimum length is increased to 60, where the model generates a summary that is not even close to the input text.
@@ 5.2.1 Function Body Generation strategy
    A set of (dynamical systems) equation implementation templates to choose from, with an informal description of the system, the states of the equations, and the necessary parameters for the equation. An example of the information required is show in Table 10, with an example implementation template written in Jinja2 (a templating language) shown in LABEL:lst:example_dynamical_equation_template. This set of templates has to be manually compiled by engineers with domain knowledge on the system of interest.
@@ 5.3 Computational model generation
When having a working code that models each of the separate components of the dynamical system, a computational model is obtained by combining these components to produce a simulation of the system that is necessary to draw conclusions about the performance of the system. This can be done by creating a highest level system class that initialises the class instances of all the components of the system, sets all the initial conditions and simulates each of them at each timestep. An example system file can be seen in LABEL:lst:example_system_class. The simulation results (i.e., the values of the states of the dynamical system) can then be used to compute performance metrics for the system or to verify whether the states exceed safe values for a system’s operation.
The generation of the system class is done by initialising all of the top-level components that subsequently initialise all their child components in an initialization function (the function called during the construction of a class instance). A class function named “simulate” is then created for the system class. This “simulate” function takes in simulation parameters and initial conditions as user input. Simulation parameters would include the length of each timestep and the total time of the simulation. The required initial conditions for user input depend on the dynamical equation being simulated. The outputs of these simulate functions will be the states of each component in the dynamical system at each timestep. These are be compiled into arrays of states at each timestep forming the output of the simulation.
@@ 6.3 Computational model generation
7 Limitations of the proposed work and possible extensions
The results presented demonstrate that the proposed approach is capable of automatically generating the computational model of a dynamical system directly from unstructured natural language text by exploiting System modeling Language (SysML) diagrams and by making strategic use of NLP strategies and very limited use of LLMs. By investigating BDD diagram results for different input texts, it was shown the advantage of the proposed approach over zero-shot LLM prompting. It was demonstrated that the generated BDD diagrams can sufficiently accurately represent the structure and components of dynamical systems and their relations. Furthermore, new validation strategies for blocks with attributes were proposed, providing a foundation for future research in this area and enabling the engineer to identify steps which might require modifications.